OpenStudy (anonymous):

f(x)=e^sqrt(4x+4) find f'(x). the answer is 4e/2sqrt(4x+4), right?

sam (.sam.):

2(4x+4)e^((4x+4)^1/2)

OpenStudy (anonymous):

can you show me the steps u used then?

OpenStudy (anonymous):

\[(e ^{\sqrt{4x+4}})' = e ^{\sqrt{4x+4}} * (\sqrt{4x+4})' * (4x+4)'\]

sam (.sam.):

Bring down the exponent, (4x+4)^1/2 e^(4x+4)^1/2 differentiate the exponent that you just bring down 2(4x+4)^3/2 e^(4x+4)^1/2

sam (.sam.):

I noticed that there is a slight error in my post, 2(4x+4) should be 2(4x+4)^-1/2

OpenStudy (anonymous):

but thats not the final answer Niker, is it?

OpenStudy (anonymous):

http://www.ms.uky.edu/~ma123/ I dont know if you can view this, but example 3 is just like this problem...

OpenStudy (anonymous):

\[2e ^{\sqrt{4x+4}} / \sqrt{4x+4}\] Sorry, uh the first time I posted I put a 4x, when it should have been just 4

OpenStudy (anonymous):

you have to click on lecture notes then on chapter 6 notes with answers

OpenStudy (anonymous):

and then?

OpenStudy (anonymous):

lecture notes ch 6 notes w/ answers example 3

OpenStudy (anonymous):

Oh, you meant suplemntary

OpenStudy (anonymous):

The answer you gave me is wrong by the way

OpenStudy (anonymous):

Why is it wrong

OpenStudy (anonymous):

o yeah, i did mean that.

OpenStudy (anonymous):

my hw website said so

OpenStudy (anonymous):

unless im putting it in wrong. its an old school site. you have to use parenthesis, ^, ^1/2, and etc...

OpenStudy (anonymous):

Tell me if it works out =/

OpenStudy (anonymous):

its not

OpenStudy (anonymous):

any suggestions?

OpenStudy (anonymous):

thats a whole different answer than what u gave me

OpenStudy (anonymous):

Actually its not, they just factored out the fours and canceled out

OpenStudy (anonymous):

this is what i put in my hw site: (2*e)^(4x+4)^(1/2))/((4x+4)^(1/2))

OpenStudy (anonymous):

(2*e)^ ( (4x+4)^(1/2)) / ((4x+4) ^(1/2))

OpenStudy (anonymous):

I dont know

OpenStudy (anonymous):

somethings wrong

OpenStudy (anonymous):

and i did have 2 parenthesis before the 4x

OpenStudy (anonymous):

((2*e)^ ( (4x+4)^(1/2)) )/ ((4x+4) ^(1/2)) Maybe add another one?

OpenStudy (anonymous):

nope...

OpenStudy (anonymous):

Does it want the super simplified form? >> Sorry if im not any help at all

OpenStudy (anonymous):

i think the answer is wrong because it doesnt say its a syntax error when i enter it

OpenStudy (anonymous):

e^[ (4x+4)^0.5] = e^2(x+1)^0.5 f'(x) = e^2[(x+1^0.5] * (x+1)^(-0.5) = e^2(x+1)^0.5 / (x+1)^(0.5)

OpenStudy (anonymous):

thats not right either :(

OpenStudy (anonymous):

ok i added some parenthesis and now its correct!!!

OpenStudy (anonymous):

ok - where are the parantheses?

OpenStudy (anonymous):

e^(2(x+1)^(1/2))/((x+1)^(1/2))

OpenStudy (anonymous):

yes ok - putting those in avoids any possible ambiguity

OpenStudy (anonymous):

i dont understand how u got 2(x+1) though

OpenStudy (anonymous):

could u explain jimmy

OpenStudy (anonymous):

Its the same thing as what i said, he just took out the 4..

OpenStudy (anonymous):

do u mean on the first line?

OpenStudy (anonymous):

yeah... e^2(x+1)^1/2

OpenStudy (anonymous):

yea sqrt (4x+4) = sqrt4 * sqrt(x+ 1) = 2(x+1)^(1/2)

OpenStudy (anonymous):

i see...